# Massive Dirac-Pauli physics in lead-halide perovskites

### v1

@@ Abstract
In standard quantum electrodynamics (QED), the so-called non-minimal (Pauli) coupling is suppressed for elementary particles and has no physical implications. Here, we show that the Pauli term naturally appears in a known family of Dirac materials – the lead-halide perovskites, suggesting a novel playground for the study of analogue QED effects. We outline measurable manifestations of the Pauli term in the phenomena pertaining to (i) the Klein paradox and (ii) relativistic corrections to bound states. In particular, we demonstrate that the binding energy of an electron in the vicinity of a positively charged defect is noticeably decreased due to the polarizability of lead ions and the appearance of a Darwin-like term. Our study adds to understanding of quantum phenomena in lead-halide perovskites, and paves the way for tabletop simulations of analogue Dirac-Pauli equations.
I Introduction
@@
Here, we show that the Pauli term naturally appears in the low-energy description of lead-halide perovskites (LHPs), making it (to the best of our knowledge) the first known Dirac-Pauli semiconductor. Lead-halide perovskites is a family of advanced energy materials [15] with exceptional optoelectronic properties [16, 17]. Charge carriers in LHPs can be described by the massive Dirac equation [18, 19] with a peculiarity that their coupling to electromagnetic fields requires an explicit inclusion of the high polarizability of Pb2+ ions [20, 21]. Here, we first illustrate that this non-standard matter-field interaction is actually the Pauli coupling with the (effective) atomic dipole polarizability in place of the anomalous moment, $\mu^{\prime}$. Further, to facilitate future theoretical and experimental studies, we outline a few effects that the Dirac-Pauli physics can have on the properties of LHPs. In particular, we consider scattering off a step potential and binding of charge carriers to local defects, such as vacancies and interstitials.
II Results
@@
where $\Psi$ is the 4-component Dirac spinor, $\Psi^{T}=(\psi_{\mathcal{\Uparrow}\uparrow},\psi_{\mathcal{\Uparrow}\downarrow},\psi_{\mathcal{\Downarrow}\uparrow},\psi_{\mathcal{\Downarrow}\downarrow})$, and $\mathcal{E}$ is the energy. [Here, the first (last) two elements of the spinor are typically referred to as electron (hole) components.] Specifically, we solve Eq. (5) with the Dirac-Pauli Hamiltonian for two standard physical problems: scattering of free charge carriers off a potential step (Klein paradox; see Fig. 1(a)) and determination of the energy of bound states in the vicinity of an attractive potential.
Klein problem in LHPs. Klein paradox refers to a counter-intuitive behavior of scattering amplitudes off a step potential whose height is larger than twice the particle’s rest mass. In the first quantization picture, the paradox manifests itself in the seemingly unphysical values for the reflection and transmission coefficients. In particular, the Klein’s solution suggests that the transmitted current does not vanish even when the height of the barrier approaches infinity [26, 27] (for more details see Refs. [28, 29, 30]).
Figure 1: (a) Schematic of the Klein problem: Electrons of energy $\mathcal{E}$ and momentum $k$ are incident on a potential barrier of height $V_{0}$. If the value of $V_{0}$ is larger than the energy gap, $\Delta$, between the conduction and valence bands, then Klein paradox occurs – particles with momentum $K$ appear in the high potential region. (b) Reflection coefficient as a function of the (dimensionless) barrier height with and without Pauli coupling assuming that $\mathcal{E}=0.01\,\text{eV}$. The other parameters correspond to MAPbBr3 with $\mu<0$. For $\mu>0$, only the sign of $\mu$ is changed in comparison to the case with $\mu<0$.
For simplicity, we assume that $t_{3}=0$ here. Similar to the original formulation of the problem by Klein, we consider the one-dimensional potential
$\phi(x)=V_{0}\Theta(x),$ (6)
where $\Theta(x)$ is the Heaviside step function, and $V_{0}$ is the strength of the potential. For $\mu\neq 0$, the associated electric field $E_{x}=-\partial\phi/\partial x=-V_{0}\delta(x)$ enters the Dirac equation via the Pauli term. This leads to a term with a $\delta$-function in Eq. (4), giving rise to a discontinuity in the spinor components at $x=0$, which can be treated as a specific boundary condition.
We solve the Dirac-Pauli equation for each spinor component and arrive at the following expressions for the reflection and transmission coefficients (see Suppl. Note 1):
where $\lambda=4iatk/(\Delta+2\mathcal{E})$ and $\Lambda=4iatK/(\Delta+2\mathcal{E}-2qV_{0})$, $r=\lambda e^{-\frac{\mu V_{0}}{at}}/\Lambda$ and $k$ and $K$ are the momenta of the charge carrier for $x<0$ and $x>0$. For $\mu\to 0$, these expressions turn into the original Klein solution [30]. Non-vanishing values of $\mu$ lead an exponential dependence on $V_{0}$ in Eq. (7). Depending on the sign of $\mu$, the exponent $e^{-\frac{\mu V_{0}}{at}}$ that appears in the parameter $r$ is either increasing or decreasing, leaving marked imprints on the transmission and reflection coefficients.
When $\mu<0$ as in LHPs, the Klein paradox is modified in comparison to the case when there is no Pauli coupling. Curiously, the introduction of the Pauli term suppresses the anomalous ($R>1$) Klein reflectivity in the asymptotic limit of large barrier heights, $qV_{0}\rightarrow\infty$. Figure 1(b) illustrates this for the parameters of MAPbBr3. The behavior of the reflection coefficient for $\mu>0$ is even more intriguing as it features a resonance-like divergent behavior, which can be interpreted as a resonant enhancement of particle-antiparicle creation within the potential barrier. This might be indicative of a novel vacuum breakdown channel that has previously been overlooked (see Suppl. Note 2 for additional details). We conclude that for the either sign of $\mu$, Pauli coupling strongly affects Klein’s physics.
@@
Bound states near static impurities in LHPs. In this subsection, we investigate how the bound states of an electron in the vicinity of a spherically symmetric local impurity potential, $\phi$, are modified by the Pauli term. To this end, we focus on the effective behavior of electrons by integrating out the hole degrees of freedom following the procedure of Ref. [35] (see Suppl. Note 3), which resembles Foldy-Wouthuysen transformation in QED [36].
@@
By doing so, we frame the problem in the form of the Schrödinger equation for electrons
$\bigg{(}-\frac{\nabla^{2}}{2m_{e}^{*}}+q\phi(r)-\frac{q^{2}\phi(r)^{2}}{\Delta}+H_{P}\bigg{)}\psi(r)=\epsilon\psi(r),$ (8)
@@
where $\psi^{T}=(\psi_{\mathcal{\Uparrow}\uparrow},\psi_{\mathcal{\Uparrow}\downarrow})$ is the electronic wavefunction; $\epsilon=\mathcal{E}-\Delta/2$ is the corresponding energy; $m_{e}^{*}=(8a^{2}t^{2}/\Delta+t_{3}a^{2}/2)^{-1}$ is the effective mass, note that the parameter $t_{3}$ enters only in the effective mass, otherwise its presence does not affect our results; $q\phi(r)-\frac{q^{2}\phi(r)^{2}}{\Delta}$ is the potential energy of an electron in the electric field generated by the impurity; $H_{P}$ is the sum of terms arising from Pauli coupling:
@@
Figure 2: (a) Ground state wavefunction with and without the Pauli term for $\frac{\mu}{aq}=0.3$. (b) Relative change in the binding energy of the electron due to the Pauli coupling plotted as a function of the binding energy with $H_{P}=0$. Note that the Pauli coupling leads to a decrease in the binding energy. Different curves are for different values of the effective dipole moment, $\mu$; the curve in the middle is calculated with the parameters of MAPbBr3.
@@
where $\varepsilon$ is the (dynamic) dielectric constant of the medium and $\mathrm{erf}$ is the standard error function. Once the electric field is known, we solve the Schrödinger equation (30) numerically by diagonalizing the corresponding matrix (see Suppl. Note 3). Specifically, we calculate the ground state energy of the system with and without $H_{P}$, and study the difference introduced by $H_{P}$.
For the ground state $L=0$, hence, the spin-orbit coupling term does not play a role in Eq. (31) and can be ignored. In other words, effects of SOC can only be seen in excited states. A study of the ground state allows us to demonstrate the effect of the first two terms in $H_{P}$: the polarizability and the Darwin terms. Note that in LHPs, the parameter $\mu$ is negative, which implies that the matrix $H_{P}$ is positive-definite, and hence, that the ground state energy is increased by the presence of the Pauli coupling.
We show our findings in Fig. 2. First, we plot the electronic density in Fig. 2(a). We see that the wave function becomes spatially broader as the Pauli term repels the electron. This change in the electronic density affects other observables as well, for example the binding energy $\epsilon_{B}>0$ of the electron-defect system. Indeed, Fig. 2(b) confirms that the Pauli term leads to a noticeable reduction of $\epsilon_{B}$. In the figure, we plot the relative decrease in binding energy:
@@
III Discussion

### v2

@@ Abstract
In standard quantum electrodynamics (QED), the so-called non-minimal (Pauli) coupling is suppressed for elementary particles and has no physical implications. Here, we show that the Pauli term naturally appears in a known family of Dirac materials – the lead-halide perovskites, suggesting a novel playground for the study of analogue QED effects. We outline measurable manifestations of the Pauli term in the phenomena pertaining to (i) relativistic corrections to bound states (ii) the Klein paradox, and (iii) spin effects in scattering. In particular, we demonstrate that (a) the binding energy of an electron in the vicinity of a positively charged defect is noticeably decreased due to the polarizability of lead ions and the appearance of a Darwin-like term, (b) strong spin-orbit coupling due to the Pauli term affects the exciton states, and (c) scattering of an electron off an energy barrier with broken mirror symmetry produces spin polarization in the outgoing current. Our study adds to understanding of quantum phenomena in lead-halide perovskites, and paves the way for tabletop simulations of analogue Dirac-Pauli equations.
@@
In this paper, we show that the Pauli term naturally appears in the low-energy description of lead-halide perovskites (LHPs), making it (to the best of our knowledge) the first known Dirac-Pauli semiconductor. Lead-halide perovskites are a family of advanced energy materials [15] with exceptional optoelectronic properties [16, 17]. Charge carriers in LHPs can be described by the massive Dirac equation [18, 19] with a peculiarity that their coupling to electromagnetic fields requires an explicit inclusion of the high polarizability of Pb2+ ions [20, 21].
Results
Here, we first illustrate that the non-standard matter-field interaction in LHPs is actually the Pauli coupling with the (effective) atomic dipole polarizability in place of the anomalous moment, $\mu^{\prime}$. Further, to facilitate future theoretical and experimental studies, we outline a few effects that the Dirac-Pauli physics can have on the properties of LHPs. We consider bound states of charge carriers to local defects, such as vacancies and interstitials, and also electron-hole binding. Finally, we discuss physics of scattering in presence of the Pauli term.
Pauli coupling in LHPs
@@ Results
where $\Psi$ is the 4-component Dirac spinor, $\Psi^{T}=(\psi_{\mathcal{\Uparrow}\uparrow},\psi_{\mathcal{\Uparrow}\downarrow},\psi_{\mathcal{\Downarrow}\uparrow},\psi_{\mathcal{\Downarrow}\downarrow})$, and $\mathcal{E}$ is the energy. [Here, the first (last) two elements of the spinor are typically referred to as electron (hole) components.] Specifically, we solve Eq. (5) with the Dirac-Pauli Hamiltonian for two standard physical problems: bound states and scattering. Specifically, we discuss binding of a charge carrier by an attractive potential and scattering of free charge carriers off a potential step (Klein paradox; see Fig. 2(a)). To guide experimental studies of the Pauli term, we further discuss spin effects in scattering and formation of excitons. The coupling between spin and the electric field is a hallmark feature of the Pauli coupling, providing a clear pathway for its detection in a laboratory.
Bound states near static impurities in LHPs
@@
In this subsection, we investigate how the bound states of an electron in the vicinity of a spherically symmetric local impurity potential, $\phi$, are modified by the Pauli term. To this end, we focus on the effective behavior of electrons by integrating out the hole degrees of freedom following the procedure of Ref. [26] (see Supplementary Note 1). This approach resembles Foldy-Wouthuysen transformation in QED [27], which is a perturbative study with the expansion parameter $\epsilon/\Delta$ ($\epsilon$ is the energy of the bound state). Perturbative approach to bound states is natural in LHPs due to its ‘defect tolerance’, i.e., only shallow states are observed in the badgap [28].
@@
By doing so, we frame the problem in the form of the Schrödinger equation for electrons that in the leading (in the powers of $1/\Delta$) order reads
$\bigg{(}-\frac{\nabla^{2}}{2m_{e}^{*}}+q\phi(r)+V_{0}+H_{P}\bigg{)}\psi(r)=\epsilon\psi(r),$ (6)
@@
where $\psi^{T}=(\psi_{\mathcal{\Uparrow}\uparrow},\psi_{\mathcal{\Uparrow}\downarrow})$ is the electronic wavefunction; $\epsilon=\mathcal{E}-\Delta/2$ is the corresponding energy; $m_{e}^{*}=(8a^{2}t^{2}/\Delta+t_{3}a^{2}/2)^{-1}$ is the effective mass, note that the parameter $t_{3}$ enters only in the effective mass, otherwise its presence does not affect our results. The terms $V_{0}$ and $H_{P}$ originate from the coupling between the conduction and valence bands. $V_{0}=-\frac{q^{2}\phi(r)^{2}}{\Delta}+\frac{2\epsilon}{\Delta}q\phi(r)$ is the beyond-Coulomb potential energy of an electron in the electric field generated by the impurity for $\mu=0$ [26]; $H_{P}$ is the sum of terms arising from Pauli coupling:
@@
Figure 1: Bound state problem in presence of Pauli coupling. (a) Ground state wavefunction with and without the Pauli term for $\frac{\mu}{aq}=0.3$ assuming the electric field from Eq. (9) with $Q=2|q|$. (b) Relative change in the binding energy of the electron due to the Pauli coupling plotted as a function of the binding energy with $H_{P}=0$. Note that the Pauli coupling leads to a decrease in the binding energy. Different curves are for different values of the effective dipole moment, $\mu$; the curve in the middle is calculated with the parameters of MAPbBr3. The marker corresponds to the parameters used in panel (a).
@@
where $\varepsilon$ is the (dynamic) dielectric constant of the medium (we take $\varepsilon=5$ in our numerical calculations [29]) and $\mathrm{erf}$ is the standard error function. Once the electric field is known, we solve the Schrödinger equation (6) numerically by diagonalizing the corresponding matrix (see Methods). Specifically, we calculate the ground state energy of the system with and without $H_{P}$, and study the difference introduced by $H_{P}$.
The ground state of Eq. (6) has $L=0$, hence, the spin-orbit coupling term does not play a role in Eq. (7) and can be ignored. In other words, effects of SOC can only be seen in the states with $L\neq 0$. A study of the ground state allows us to demonstrate the effect of the first two terms in $H_{P}$: the polarizability and the Darwin terms. Note that in LHPs, the parameter $\mu$ is negative, which implies that the matrix $H_{P}$ is positive-definite, and hence, that the ground state energy is increased by the presence of the Pauli coupling.
We show our findings in Fig. 1. First, we plot the electronic density in Fig. 1(a). We see that the wave function becomes spatially broader as the Pauli term repels the electron. This change in the electronic density affects other observables as well, for example the binding energy $\epsilon_{B}>0$ ($\epsilon_{B}=-\epsilon$) of the electron-defect system. Indeed, Fig. 1(b) confirms that the Pauli term leads to a noticeable reduction of $\epsilon_{B}$ (see also Supplementary Figure 1). In the figure, we plot the relative decrease in binding energy:
@@
Spin structure of excitons in LHPs
LHPs are known for their anomalously fast photon emission at cryogenic as well as at room temperatures [30, 19]. It has been suggested that this is because excitons in LHPs have a unique fine structure with an optically active (“bright”) ground state. Such level ordering can be obtained by introducing a Rashba-type spin-orbit interaction [19, 31]. In general, there is a consensus that Rashba coupling is featured prominently in LHPs, however it is usually introduced phenomenologically to account for experimental observations. The microscopic origin of Rashba coupling is a subject of debate, especially at high temperatures where LHPs have a nominally cubic structure [32, 33, 34].
With the Pauli term, Rashba coupling follows naturally, with a coefficent determined by the parameters of the perovskite structure (cf. Ref. [21]):
$H_{\mathrm{Rashba}}=-\frac{4at\mu}{\Delta}\tau_{3}\otimes\mathbf{E}\cdot(\mathbf{k}\times\bm{\sigma}),$ (11)
where $\mathbf{E}$ is the electric field explicitly breaking inversion symmetry. The Rashba energy that corresponds to $H_{\mathrm{Rashba}}$ is
$E_{\mathrm{Rashba}}\simeq\frac{8a^{2}t^{2}\mu^{2}\langle\mathbf{E}^{2}\rangle}{\Delta^{2}}m_{e}^{*}.$ (12)
According to Ref. [19], the lowering of the “bright” state is given by $\delta E_{X}\simeq-E_{\mathrm{Rashba}}$. To compensate for the standard singlet-triplet splitting ($\simeq 0.3$meV [19]) due to the electron-hole exchange interaction [35], one should have $\delta E_{X}\simeq 1$meV. Under the hypothesis of local polar fluctuations [36], this implies the fields $|\mathbf{E}|\sim 0.6$V/nm.
While internal fields of this strength are reasonable on the atomic scales [37], they are quite considerable and are yet to be experimentally confirmed. This requirement for the strength of these fields is high because of the factor $1/\Delta^{2}$ in Eq. (12), which follows from the use of second-order perturbation theory. Surprisingly, it turns out that Pauli term can lower the optically active exciton state at a lower order of perturbation theory, thus producing the desired level splitting with lower electric fields ($|\mathbf{E}|\sim 0.3$V/nm). Indeed, the effective polarizability of a charge carrier due to Pauli coupling implies a van-der-Waals-like lowering of the energy by
$\delta E_{X}\simeq-\frac{\mu^{2}\mathbf{E}^{2}}{\Delta},$ (13)
which enters in the first order in $1/\Delta$, in contrast to the result in Eq. (12).
We conclude that the Pauli coupling may play a role in the observed fast photon emission in LHPs. Note that all derivations in this work apply to a cubic phase of LHPs. However, the polarizability of lead ions is a local property of the material, thus, we expect that our discussion applies also at lower temperatures, where LHPs crystals are tetragonal or orthorhombic.
Klein problem in LHPs
Klein paradox refers to a counterintuitive behavior of scattering amplitudes off a step potential whose height is larger than twice the particle’s rest mass. In the first quantization picture, the paradox manifests itself in the seemingly unphysical values for the reflection and transmission coefficients. In particular, the Klein’s solution suggests that the transmitted current does not vanish even when the height of the barrier approaches infinity [38, 39] (for more details see Refs. [40, 41, 42]).
Figure 2: Klein problem in LHPs. (a) Schematic of the Klein problem: Electrons of energy $\mathcal{E}$ and momentum $k$ are incident on a potential barrier of height $V_{0}$ (recall that the associated electric field is $E_{x}=-V_{0}\delta(x)$). If the value of $V_{0}$ is larger than the energy gap, $\Delta$, between the conduction and valence bands, then Klein paradox occurs – particles with momentum $K$ appear in the high potential region. (b) Reflection coefficient as a function of the (dimensionless) barrier height with and without Pauli coupling assuming that $\mathcal{E}=\Delta/2+0.01\,\text{eV}$. The other parameters correspond to MAPbBr3 with $\mu<0$. For $\mu>0$, only the sign of $\mu$ is changed in comparison to the case with $\mu<0$.
For simplicity, we assume that $t_{3}=0$ here  [^3] Otherwise, we should explicitly consider other electronic bands whose presence leads to $t_{3}\neq 0$. Similar to the original formulation of the problem by Klein, we consider the one-dimensional potential
$\phi(x)=V_{0}\Theta(x),$ (14)
where $\Theta(x)$ is the Heaviside step function, and $V_{0}$ is the strength of the potential. For $\mu\neq 0$, the associated electric field $E_{x}=-\partial\phi/\partial x=-V_{0}\delta(x)$ enters the Dirac equation via the Pauli term. This leads to a term with a $\delta$-function in Eq. (4), giving rise to a discontinuity in the spinor components at $x=0$, which can be treated as a specific boundary condition, see Supplementary Note 2 and Ref. [48].
We solve the Dirac-Pauli equation for each spinor component and arrive at the following expressions for the reflection and transmission coefficients (see Supplementary Note 2):
where $\lambda=4iatk/(\Delta+2\mathcal{E})$ and $\Lambda=4iatK/(\Delta+2\mathcal{E}-2qV_{0})$, $r=\lambda e^{-\frac{\mu V_{0}}{at}}/\Lambda$ and $k$ and $K$ are the momenta of the charge carrier for $x<0$ and $x>0$. For $\mu\to 0$, these expressions turn into the original Klein solution [42]. Non-vanishing values of $\mu$ lead to an exponential dependence on $V_{0}$ in Eq. (15). Subject to the sign of $\mu$, the exponent $e^{-\frac{\mu V_{0}}{at}}$ that appears in the parameter $r$ is either increasing or decreasing, leaving marked imprints on the transmission and reflection coefficients.
When $\mu<0$ as in LHPs, the Klein paradox is modified in comparison to the case when there is no Pauli coupling. Curiously, the introduction of the Pauli term suppresses the anomalous ($R>1$) Klein reflectivity in the asymptotic limit of large barrier heights, $qV_{0}\rightarrow\infty$. Figure 2(b) illustrates this for the parameters of MAPbBr3. The behavior of the reflection coefficient for $\mu>0$ is even more intriguing as it features a resonance-like divergent behavior, which can be interpreted as a resonant enhancement of particle-antiparicle creation within the potential barrier. This might be indicative of a novel vacuum breakdown channel that has previously been overlooked (see Supplementary Note 3 for additional details). We conclude that for the either sign of $\mu$, Pauli coupling strongly affects Klein’s physics.
Spin effects in scattering
Without doubt, modifications to Klein’s physics is of general interest, however, their observation appears challenging due to the energy gap in LHPs. As we discuss below, the effects of Pauli coupling on scattering are more readily studied through the spin degree of freedom. To illustrate this, let us assume that a sample of LHPs has an energy barrier described by the potential $\phi(x)=v_{0}f(x)$, where $v_{0}$ ($f$) is the strength (shape) of the barrier. By assumption, the function $f(x)$ has a finite support parametrized by $d$, i.e., $f(x)=0$ if $|x|>d/2$. One realization of the potential $\phi(x)$ is an n-p-n junction [49]. Unlike the Klein scenario, the potential barrier here is weak, $v_{0}\ll\Delta$, allowing us to implement Foldy-Wouthuysen-like transformation and derive the equation for electron propagation, cf. Eq. (6):
$\bigg{(}-\frac{\nabla^{2}}{2m_{e}^{*}}+W(x)+\frac{4at\mu v_{0}}{\Delta}\frac{\mathrm{d}f}{\mathrm{d}x}\sigma_{z}k_{y}\bigg{)}\psi(\mathbf{r})=\epsilon\psi(\mathbf{r}),$ (16)
where we chose the $z$-axis perpendicular to the scattering plane; $W(x)$ is the spin-independent part of the potential (see Supplementary Note 1); $k_{y}$ is the quantum number that parametrizes the momentum along the $y$-direction.
The Pauli coupling in Eq. (16) turns the energy barrier into a magnetic impurity, which affects spin-up and spin-down electrons differently, allowing one to manipulate the spin of an electron. For example, given a spin-unpolarized current for $x<0$, one can generate a spin-polarized current for $x>d$. Indeed, let us consider the transmission coefficient (see, e.g., Ref. [50])
$T_{\pm}=\left|1+\frac{im_{e}^{*}}{k_{x}}\int\mathrm{d}xW_{\pm}(x)e^{-ik_{x}x}\psi_{1}(x)\right|^{-2}.$ (17)
Here, $\pm$ corresponds to the spin of the incoming electron; $W_{\pm}=W\pm 4atv_{0}\mu f^{\prime}k_{y}/\Delta$. The function $\psi_{1}$ is the solution to Eq. (16) that satisfies the integral equation
$\psi_{1}(x)=e^{ik_{x}x}-\frac{2m_{e}^{*}}{k_{x}}\int_{x}^{\infty}\mathrm{d}y\sin[k_{x}(x-y)]W_{\pm}(y)\psi_{1}(y),$ (18)
where the momentum $k_{x}$ parametrizes the incoming flux. As $W_{+}\neq W_{-}$, we conclude that $T_{+}\neq T_{-}$, which leads to the spin polarization of the outgoing current: $|T_{+}-T_{-}|/(T_{+}+T_{-})$. To measure it in LHPs, one can use the standard methodology established for other semiconductors [51].
The simplest estimate for $T_{\pm}$ is obtained in the quantum tunneling regime using the WKB approximation, $T^{\mathrm{WKB}}_{\pm}\sim\exp(-2\int_{x_{1}}^{x_{2}}\sqrt{2m_{e}^{*}(W_{\pm}-\epsilon)})$, where $x_{1}$ and $x_{2}$ are classical turning points. This expression allows one to optimize $f(x)$ for the problem at hand. It is worth noting that $T^{\mathrm{WKB}}_{+}\neq T^{\mathrm{WKB}}_{-}$ only for potentials that break mirror symmetry, i.e., if $f(-x)\neq f(x)$. This is in line with the recent observation of spin polarized currents in chiral LHPs at room temperature in the absence of external magnetic fields [52]. Our results demonstrate that similar observations can be made in non-chiral LHPs in the presence of an energy barrier with broken mirror symmetry.
Discussion